The temperature of a gas is 298.0 k when the pressure is 4.00 atm. At the same volume, what is the temperature at which the pressure of the gas is 3.0 atm
1
Expert's answer
2018-05-23T06:27:36-0400
1 = 298 K

P1 = 4 atm

P2 = 3 atm

Gay-Lussac`s Law: P1/T1 = P2/T2

T2 = T1*P2/P1 = 298K*3atm/4atm = 223.5 K